# Rajamanickam

### v3

@@ Acknowledgment
$\mathbf{H}_{\bar{\mathbf{q}}}\mathbf{q}=\sigma\mathbf{q}$ (79)
@@
$\mathbf{H}_{\bar{\mathbf{q}}}=\left.\frac{\delta^{2}F}{\delta q_{1}\delta q_{2}}\right|_{\bar{\mathbf{q}}}=2\begin{bmatrix}-\nabla^{2}&0\\ 0&-\nabla^{2}\end{bmatrix}+2\varepsilon^{2}\begin{bmatrix}12\bar{q}_{1}^{2}+4\bar{q}_{2}^{2}-1&8\bar{q}_{1}\bar{q}_{2}\\ 8\bar{q}_{1}\bar{q}_{2}&4\bar{q}_{1}^{2}+12\bar{q}_{2}^{2}-1\end{bmatrix}$. (80)
@@
Stability in the Oseen–Frank limit
It is instructive to deduce the formulation of the stability analysis from (79) for the leading-order Oseen–Frank problem described in (35)-(36). For shortness, we shall drop the subscript ”0” for the leading-order solution. First of all, let us substitute $\bar{s}^{2}=4\bar{q}^{2}=1$, so that the eigenvalue equation reduces to
$\nabla^{2}\hat{\mathbf{q}}-8\varepsilon^{2}\bar{\mathbf{q}}\bar{\mathbf{q}}^{T}\hat{\mathbf{q}}=-\frac{\sigma}{2}\hat{\mathbf{q}}$. (82)
The matrix $\bar{\mathbf{q}}\bar{\mathbf{q}}^{T}$ is of rank one, with eigenvalues $(\frac{1}{4},0)$ and corresponding eigenvectors $(\mathbf{e}_{\parallel}=\bar{\mathbf{q}}=(\bar{q}_{1},\bar{q}_{2})^{T},\mathbf{e}_{\perp}=(-\bar{q}_{2},\bar{q}_{1})^{T})$. Hence any perturbation can be decomposed as
$\hat{\mathbf{q}}=\hat{q}_{{}_{\parallel}}\mathbf{e}_{\parallel}+\hat{q}_{\perp}\mathbf{e}_{\perp}.$ (83)
@@
However, in the Oseen–Frank limit, $\varepsilon\to\infty$, the linear perturbation must remain on the manifold $s=1$, so that only the perpendicular component is dynamically relevant. The resulting eigenvalue equation reads
@@
where the term $4|\nabla\bar{\varphi}|^{2}\hat{q}_{\perp}$ arises from the spatial variation of the basis vector $\mathbf{e}_{\perp}(\mathbf{x})$. Introducing $\hat{q}_{\perp}=2\hat{\varphi}$, with $\hat{\varphi}$ denoting the perturbation of the director field, one obtains
$\nabla^{2}\hat{\varphi}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{\varphi}=0$ (85)
subject to $\hat{\varphi}=0$ on $\partial\Omega$. This eigenvalue problem is formally analogous to the two-dimensional Schrödinger equation
@@
Since the potential $4|\nabla\bar{\varphi}|^{2}\geq 0$, it follows that the energy $\sigma/2\geq 0$, indicating that the base state is linearly stable. This reflects the strength and limitation of the Oseen–Frank approximation because elastic stability of the director field is captured while instabilities associated with variations in the magnitude of $\mathbf{Q}$ are neglected. In the full Landau–de Gennes theory, variations in $s$ can produce negative eigenvalues corresponding to core instabilities of defects. Within such cores, both $\hat{q}_{\parallel}$ and $\hat{q}_{\perp}$ must be retained, and the above analysis provides only an approximate description.

### v4

@@ Acknowledgment
$\mathbf{H}_{\bar{\mathbf{q}}}\hat{\mathbf{q}}=\sigma\hat{\mathbf{q}}$ (79)
@@
$\mathbf{H}_{\bar{\mathbf{q}}}=\left.\frac{\delta^{2}F}{\delta q_{1}\delta q_{2}}\right|_{\bar{\mathbf{q}}}=-2\mathbf{I}\nabla^{2}+2\varepsilon^{2}[8\bar{\mathbf{q}}\otimes\bar{\mathbf{q}}+(4|\bar{\mathbf{q}}|^{2}-1)\mathbf{I}].$ (80)
@@
The Eigenvalue problem in a natural intrinsic basis
To analyse stability more transparently, it is convenient to project perturbations onto a local intrinsic frame aligned with the unperturbed $\mathbf{Q}$-tensor. The matrix $\bar{\mathbf{q}}\otimes\bar{\mathbf{q}}$ is of rank one, with eigenvalues $(\frac{1}{4},0)$ and corresponding normalised eigenvectors
$\mathbf{e}_{\parallel}=\frac{\bar{\mathbf{q}}}{|\bar{\mathbf{q}}|},\qquad\mathbf{e}_{\perp}=\frac{1}{|\bar{\mathbf{q}}|}\begin{pmatrix}-\bar{q}_{2}\\ \bar{q}_{1}\end{pmatrix}$. (82)
Any perturbation can then be decomposed as
$\hat{\mathbf{q}}=\hat{q}_{{}_{\parallel}}\mathbf{e}_{\parallel}+\hat{q}_{\perp}\mathbf{e}_{\perp}$ (83)
where $\hat{q}_{\parallel}$ represents variations in the magnitude of the $\mathbf{Q}$-tensor and $\hat{q}_{\perp}$ corresponds to reorientation of the director. The above decomposition decouples the parallel and perpendicular modes in the bulk potential, though the elastic term becomes more intricate due to the spatial variation of the basis vectors.
The local rotation of the unperturbed director is captured by the one-form
$\boldsymbol{\omega}=\mathbf{e}_{\perp}\cdot\nabla\mathbf{e}_{\parallel}=\frac{\bar{q}_{1}\nabla\bar{q}_{2}-\bar{q}_{2}\nabla\bar{q}_{1}}{|\bar{\mathbf{q}}|^{2}}=2\nabla\bar{\varphi}$ (84)
so that $\nabla e_{\parallel}=\boldsymbol{\omega}\mathbf{e}_{\perp}$ and $\nabla e_{\perp}=-\boldsymbol{\omega}\mathbf{e}_{\parallel}$. Projecting the Hessian equation, $\mathbf{H}_{\bar{\mathbf{q}}}\hat{\mathbf{q}}=\sigma\hat{\mathbf{q}}$, onto the intrinsic frame yields
$\Delta_{\boldsymbol{\omega}}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}-\varepsilon^{2}\begin{bmatrix}12|\bar{\mathbf{q}}|^{2}-1&0\\ 0&4|\bar{\mathbf{q}}|^{2}-1\end{bmatrix}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}=-\frac{\sigma}{2}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}$ (85)
where $\Delta_{\boldsymbol{\omega}}$ is the covariant Laplacian, defined by
$\Delta_{\boldsymbol{\omega}}$ $=\mathbf{I}(\nabla^{2}-|\boldsymbol{\omega}|^{2})+\mathbf{J}(2\boldsymbol{\omega}\cdot\nabla+\nabla\cdot\boldsymbol{\omega}),\qquad\mathbf{J}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$, (86)
$=\mathbf{I}(\nabla^{2}-4|\nabla\bar{\varphi}|^{2})+\mathbf{J}(4\nabla\bar{\varphi}\cdot\nabla+2\nabla^{2}\bar{\varphi}).$ (87)
As we can see, the $90^{\circ}$-rotation matrix $\mathbf{J}$ couples the parallel and perpendicular modes. Although $\mathbf{J}$ is skew-symmetric, the operator $\mathbf{J}(2\boldsymbol{\omega}\cdot\nabla+\nabla\cdot\boldsymbol{\omega})$ is self-adjoint, as can be verified directly, ensuring that the eigenvalues remain real. The first term, $\mathbf{I}\nabla^{2}$, represents the standard elastic penalty and acts to smooth perturbations. The second term, $4|\nabla\bar{\varphi}|^{2}\mathbf{I}$, acts like a potential (see below), rendering perturbations more difficult to sustain in regions where the background director varies strongly. The third term, $2\mathbf{J}\boldsymbol{\omega}\cdot\nabla$, captures how perturbations in one mode (say, parallel amplitude changes) get converted to the other mode (perpendicular director angle changes) as one follows the background director rotation. The fourth term, $\mathbf{J}\nabla\cdot\boldsymbol{\omega}$, acts like a local source that converts between perturbation modes based on the local accumulation of director rotation and plays a significant role in defect cores, where $\nabla^{2}\bar{\varphi}\neq 0$ in general. Together, these four contributions fully encode the interplay between director rotation, mode coupling, and elastic smoothing in the linearised dynamics around the equilibrium state.
The bulk potential terms in the Hessian are diagonal in this basis. Since $|\bar{\mathbf{q}}|^{2}\in[0,1/4]$, we have $12|\bar{\mathbf{q}}|^{2}-1\in[-1,2]$ and $4|\bar{\mathbf{q}}|^{2}-1\in[-1,0]$. The parallel term $12|\bar{\mathbf{q}}|^{2}-1$ thus stabilises magnitude fluctuations when $\bar{s}$ is close to unity, but can destabilize in low-order regions such as defect cores. The perpendicular term $12|\bar{\mathbf{q}}|^{2}-1$ governs how director rotations couple to the local degree of order. In perfectly ordered regions, where $\bar{s}=1$, this term is zero, it vanishes, reflecting rotational symmetry and producing soft elastic modes. In regions of reduced order, it is negative, meaning the bulk potential provides no restoring force for director rotations, which then rely primarily on elasticity for stability. Together, these effects imply that defect cores are inherently vulnerable to instabilities in both magnitude and orientation, a phenomenon absent in the director‑only Oseen–Frank approximation.
Director stability in the Oseen–Frank limit
The results for the Oseen–Frank limit is simple. For brevity, consider only the leading-order Oseen–Frank problem described in (35)-(36) and drop the subscript ”0”. Since $\bar{s}^{2}=4|\bar{\mathbf{q}}|^{2}=1$ and $\nabla^{2}\bar{\varphi}=0$, the eigenvalue equation (85) simplifies to
$\Delta_{\boldsymbol{\omega}}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}-\varepsilon^{2}\begin{bmatrix}2&0\\ 0&0\end{bmatrix}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}=-\frac{\sigma}{2}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}$ (88)
@@
with $\Delta_{\boldsymbol{\omega}}=\mathbf{I}(\nabla^{2}-4|\nabla\bar{\varphi}|^{2})+4\mathbf{J}\nabla\bar{\varphi}\cdot\nabla$. However, in the Oseen–Frank limit, $\varepsilon\to\infty$, the linear perturbation must remain on the manifold $s=1$ because $\hat{q}_{\parallel}=O(\varepsilon^{-2})$, so that only the perpendicular component is dynamically relevant. The resulting eigenvalue equation reads
@@
Introducing $\hat{q}_{\perp}=2\hat{\varphi}$, with $\hat{\varphi}$ denoting the perturbation of the director field, one obtains
$\nabla^{2}\hat{\varphi}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{\varphi}=0,\qquad\hat{\varphi}_{\partial\Omega}=0.$ (90)
This eigenvalue problem is formally analogous to the two-dimensional Schrödinger equation
@@
Since the ”potential” $4|\nabla\bar{\varphi}|^{2}\geq 0$, it follows that the ”energy” $\sigma/2\geq 0$, indicating that the base state is linearly stable. This reflects the strength and limitation of the Oseen–Frank approximation because elastic stability of the director field is captured while instabilities associated with variations in the magnitude of $\mathbf{Q}$ are neglected. In the full Landau–de Gennes theory, variations in $s$ can produce negative eigenvalues corresponding to core instabilities of defects. Within such cores, both $\hat{q}_{\parallel}$ and $\hat{q}_{\perp}$ must be retained, and the above analysis provides only an approximate description.